User Tools

Site Tools


handover:r1703

R1703

Ho

VSN: USN-0111

  • UT 1701 experiment started OK. Can't find r1703ho.sum, prc files etc seem to have drudged okay, but oper@hobart /tmp/sched.tmp looks to be for previous experiment still. (JS)
  • UT1730 The default attenuations are sufficient to keep the voltages in the appropriate range - leaving them (JS)
  • UT0650 vacuum pump switched on because receiver was slowly warming up. (Lucia)

Ke

VSN: HOB+0047 0GB

  • UT 1704 experiment started okay (JS)

Yg

VSN: USN-0035 0GB

  • UT 1704 experiment started okay (JS)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1703.txt · Last modified: 2015/09/01 06:50 by Lucia Plank